OpenStudy (anonymous):

x+y+z=3 x-y+2z=9 5x+y+z=15

OpenStudy (athe):

x=3-y-z -2y+z=6 -4y-4z=0 y=-z x=3 z=2y+6 x=3 z=-y z=-2z+6 x=3 z=2 y=-2